# Towards Principled Superhuman AI for Multiplayer Symmetric Games

### v1

@@ 3 Equilibria and Self-play are Insufficient for Multiplayer Games
In each round, every player chooses either 00 or $1$, with the rule that being in the majority yields a positive payoff of $1$, while being in the minority results in a negative payoff of $-2$.
Limitation of equilibria
@@
In this setup, both strategies $(0,0,0)$ and $(1,1,1)$ constitute NE. However, the existence of multiple NEs creates a predicament for the AI agent. It must choose which equilibrium to follow, yet there’s always the risk that the other two players may coordinate on the opposing NE, leading to a negative payoff for the AI agent. In other words, adhering to a single NE does not reliably ensure an average payoff when multiple equilibria exist. Since ${\rm NE}\subset{\rm CE}\subset{\rm CCE}$, we know the same limitation also holds for CE and CCE.
Limitation of self-play from scratch
Self-play is a training method in which an AI agent improves its performance by repeatedly playing against copies of itself without human supervision. Typical self-play is carried out by optimizing one agent’s policy (denoted as the main agent) while fixing all other agents’ policies to be the current policy or the earlier versions of the main agent. When no-regret learning algorithms such as Hedge are adopted as the optimizer, the self-play algorithm is guaranteed to converge to a CCE. Moreover, it may converge to a certain “good” subclass of CCEs depending on the choice of specific optimizer and initialization.
Here, we argue that self-play from scratch (Brown and Sandholm,, 2019) fails to always achieve non-negative expected payoff in multiplayer symmetric zero-sum games. The example is again the three-player majority vote game. Imagine two parallel worlds, with two other players always playing 1 in one world, and 0 in the other. The learner must play the exact same action as the other two players to receive non-negative expected payoff. Noting that self-play from scratch does not check other players’ actual policy, the learner has no clue which world it is, and thus loses in at least one of the two worlds.
We note that several recent systems (Li et al.,, 2020; Jacob et al.,, 2022) combine self-play with human policy modeling which avoids the drawback of self-play from scratch. However, our experiments demonstrate that many of these solutions remain ineffective even in some simple normal-form games (See Section 6).
4 New Solution Concepts
The limitation of standard equilibria leads us to the following question: What solution concept should be adopted to secure an equal share (non-negative expected payoff) in multiplayer symmetric games? In this section, we provide a clear answer to this question.
Diverse strategies by opponents
We start by considering the general cases where opponents can possibly deploy different strategies.
Claim 4.1.
There exist symmetric zero-sum games such that no policy achieves non-negative expected payoff if the other players, even without collusion, are permitted to adopt different strategies.
Here, we define “no collusion” by requiring the policies of other players to be statistically independent without any shared randomness.
Proof.
We consider a $3$-player minority game involving an AI agent and two other players. The rule of the game is that being in the minority yields a positive payoff of $2$, while being in the majority results in a negative payoff of $-1$. If the other two players act 00 and $1$, respectively, then the AI agent can only receive $-1$ payoff, which is strictly less than 00. Since the other two players are simply playing deterministic policies, they are independent, i.e., without collusion. ∎
This observation underscores a crucial aspect of the multiplayer symmetric game dynamic where diverse strategies by opponents can drastically affect the possibility of achieving an equal share.
Identical strategy by opponents
Claim 4.1 leaves out no choice but to consider the setting where opponents must deploy identical strategy. While it may seem limiting to assume that opponents employ identical strategies, we argue that this assumption is actually already implicit in the majority of practical systems (Li et al.,, 2020; Brown and Sandholm,, 2019; Bakhtin et al.,, 2022; † et al., 2022, FAIR). These systems utilize a self-play framework, equating the strategies of all opponents with those of the learner. Identical strategy assumption is also naturally satisfied in modern multiplayer gaming platforms with a large player base (See Section 4.1).
After making the strategies of opponents equal, one may wonder whether we can find a fixed non-exploitable policy similar to the case of two-player zero-sum games.
Claim 4.2.
There exist symmetric zero-sum games such that no fixed strategy guarantees a non-negative expected payoff against adversarial strategies by opponents, even under the constraint that they adhere to identical strategies.
Proof.
Imagine an AI agent involves in a $3$-player majority game and plays a mixed strategy $(\beta,1-\beta)$ (i.e. play 00 w.p. $\beta$; play $1$ w.p. $1-\beta$). And the other two players adopt an identical mixed strategy $(p,1-p)$ (i.e. play 00 w.p. $p$; play $1$ w.p. $1-p$). Then, we can calculate the payoff of the AI agent as $U_{1}(\beta,p,p)=\beta(-2(1-p)^{2}+2p(1-p))+(1-\beta)(-2p^{2}+2p(1-p))$. It then follows that $\max_{\beta\in[0,1]}\min_{p\in[0,1]}U_{1}(\beta,p,p)\leq\max_{\beta\in[0,1]}\min_{p\in\{0,1\}}U_{1}(\beta,p,p)=-1,$ which is strictly less than 00. ∎
This observation again makes multiplayer games significantly different from two-player zero-sum games. This leads to the following conclusion:
To reliably secure an equal share in symmetric games, an agent must be in a scenario where: (a) opponents employ the identical strategy, and (b) the AI agent adapts its strategy to the opponents’ strategies, which involves opponent modeling.
It is not difficult to find solution concepts—such as (1) behavior cloning for the strategy by opponents; and (2) best response to the identical strategy by opponents—that satisfy both criteria, and are guaranteed to secure an equal share.
Finally, we rigorously summarize our discussion using mathematical language, detailing the relationships between various minimax quantities.
@@ 4 New Solution Concepts
For symmetric zero-sum games, we have
$\max_{x_{1}}\min_{x_{2},\cdots,x_{n}}U_{1}(x_{1},\cdots,x_{n})\leq\min_{x_{2},\cdots,x_{n}}\max_{x_{1}}U_{1}(x_{1},\cdots,x_{n})\leq$ $\min_{x}\max_{x_{1}}U_{1}(x_{1},x^{\otimes n-1})=0$
$\max_{x_{1}}\min_{x}U_{1}(x_{1},x^{\otimes n-1})\leq$ $\min_{x}\max_{x_{1}}U_{1}(x_{1},x^{\otimes n-1})=0,$
where all the inequalities can be made strict in certain games.
4.1 Multiplayer games with a large player base
We argue that assumption of all opponents playing identical strategy is well-justified in modern multiplayer gaming platforms with a large player base. Imagine a casino hosting $N$ players who randomly join poker tables, or an online Mahjong match-making platform with $N$ users. Let $\{x_{i}\}^{N}_{i=1}$ be the strategy set for these $N$ players. We can then define the meta-strategy of the population as $\bar{x}=(1/N)\sum_{i=1}^{N}x_{i}$. The following proposition claims that, for $n$-player symmetric zero-sum games, as long as $N\gg(n-2)^{2}$, the payoff achieved by playing against $n-1$ players that uniformly draw from the player pool is almost the same as playing against $n-1$ players who all adopt the same meta-strategy $\bar{x}$.
@@ 4.1 Multiplayer games with a large player base
Let $\mathbb{E}_{x_{-1}}$ be the expectation over the randomness on sampling $n-1$ strategies uniformly from the set $\{x_{i}\}^{N}_{i=1}$ without replacement. Then for any policy $z\in\Delta(\mathcal{A})$, we have $|\mathbb{E}_{x_{-1}}[U_{1}(z,x_{-1})]-U_{1}(z,\bar{x}^{\otimes n-1})|\leq 2(n-2)^{2}/N$.
5 Efficient Algorithms
In this section, we explore efficient algorithms that are able to secure an average payoff within a framework where all opponents adopt a common strategy $y^{t}$ for each round $t\in[T]$. Correspondingly, the payoff function of the AI agent is given by $u^{t}(\cdot):=U_{1}(\cdot,(y^{t})^{\otimes n-1})$. Given that the rules of the game are known to the player, the player is actually aware of $U_{1}$. We approach this exploration through two distinct scenarios: (i) Section 5.1 examines scenarios with a stationary opponent, wherein $y^{t}$ remains the same over time; (ii) Section 5.2 transitions to considering slowly changing adversarial opponents, limiting their power by imposing a variation budget.
@@ 5.1 Fixed opponents
We begin by exploring the simple stationary scenario, where the common strategy adopted by the opponent remains constant over time, denoted as $y^{t}=y$ for all $t\in[T]$.
Notably, in this particular scenario, the payoff function $u^{t}(\cdot)$ remains constant over time. Additionally, by symmetry, it is observable that $\max_{a\in\mathcal{A}}U_{1}(a,y^{\otimes n-1})\geq 0,$ indicating the existence of at least one action that can consistently yield an average payoff in each round. Thus, any no-regret algorithm is poised to achieve zero payoff under this setting. For instance, we will deploy the Hedge algorithm, which provides the following guarantees:
@@
$(1/T)\sum^{T}_{t=1}u^{t}(x^{t})\geq u^{\star}-C\sqrt{\log(A/\delta)/T},$
for some absolute constant $C>$, where $u^{\star}:=\max_{a\in\mathcal{A}}U_{1}(a,y^{\otimes n-1})\geq 0$.
Theorem 5.1 claims that with stationary opponents, the Hedge algorithm is capable of achieving approximately an equal share when $T$ is large, demonstrating its effectiveness in the long term.
@@ 5.2 Adaptive opponents
In practical scenarios, encountering a fixed opponent strategy is relatively uncommon. More often, opponents adapt and modify their strategies over time, responding to the game’s dynamics and the actions of other players. Thus, in this section, we shift our focus to the non-stationary scenario, where the common strategy $y^{t}$ adopted by the opponent varies over time.
Challenges when facing fast adapting opponents
In multiplayer games, a significant challenge arises if opponents can change their strategies arbitrarily fast as demonstrated by the following fact.
Fact 5.2.
For a $3$-player majority game, there exists a stochastic evolution such that the maximum total payoff achievable by any algorithm over the duration of the game is at most $-T$.
In light of this negative result, it is clear that attaining an average payoff is impractical if arbitrary evolution is allowed. Thus we introduce a constraint on this evolution by positing a variation budget $V_{T}$, which bounds the total variation of the payoff function across the time horizon. Specifically, we assume the payoff function belongs to $\mathcal{U}$, which is defined as
$\textstyle\mathcal{U}:=\left\{\{u^{t}\}^{T}_{t=1}\,\Bigg{|}\,\sum_{t=1}^{T-1}\|u^{t+1}-u^{t}\|_{\infty}\leq V_{T}\right\}.$ (2)
Furthermore, we denote $\mathcal{G}(n,A,V_{T})$ as the set of symmetric zero-sum games involving $n$ players, $A$ actions, and the payoff function $\{u^{t}\}^{T}_{t=1}\in\mathcal{U}$. This constraint effectively moderates the power of the opponents compared to a fully adversarial setup.
@@
Minimizing standard regret is only effective in the stationary environment when a fixed action consistently yields a satisfactory payoff. In non-stationary environments, we turn our attention to dynamic regret, defined as:
@@
Note that in the setting of symmetric games, the dynamic oracle is always assured to secure an equal share, i.e., $\sum_{t=1}^{T}\max_{a\in\mathcal{A}}u^{t}(a)\geq 0$. Thus, if an algorithm can achieve no-dynamic-regret, then it’s guaranteed to achieve zero expected payoff even as the scenario evolves over time. To this ends, we adapt a no-dynamic-regret algorithm—Strongly Adaptive Online Learner with Hedge $\mathcal{H}$ as a blackbox algorithm ($\mathrm{SAOL}^{\mathcal{H}}$), as proposed by (Daniely et al.,, 2015) to our setting and achieve following guarantees:
@@
[…]
Theorem 5.3 implies that the payoff of $\mathrm{SAOL}^{\mathcal{H}}$ achieves an average payoff $(1/T)\sum^{T}_{t=1}u^{t}(x^{t})\geq\sum_{t=1}^{T}\max_{a\in\mathcal{A}}u^{t}(a)-\tilde{O}(V_{T}^{1/3}T^{-1/3})$. Therefore, if $V_{T}$ is sublinear in $T$, $\mathrm{SAOL}^{\mathcal{H}}$ is capable of approximately achieving zero expected payoff over an extended duration.
Middle regime
Interesting, there is a regime in the middle where opponents are changing reasonably fast while the value of dynamic oracle $\sum_{t=1}^{T}\max_{a\in\mathcal{A}}u^{t}(a)$ is not high. In this case, the favorable policy for the learner might be simply behavior cloning—simply mimic opponents’ strategies.
Formally, we define the behavior cloning algorithm by the learner making her action in $t$-th round the same as the action taken by the 2nd player in $(t-1)$-th round (See Algorithm 1). Behavior cloning achieves the following:
@@
Suppose that $n\geq 3$, $A\geq 2$, and $V_{T}\in[1,T]$, then behavior cloning guarantees that $\inf_{\mathcal{G}(n,A,V_{T})}\mathbb{E}\left[\sum^{T}_{t=1}u^{t}(x^{t})\right]\geq-V_{T}-1,$ where the expectation is taken over the noisy actions.
Theorem 5.4 shows that the expected average total payoff achieved by behavior cloning is at least $O(-V_{T}/T)$. We note that this can be better than the guarantees of $\mathrm{SAOL}^{\mathcal{H}}$ when the measurement for change $V_{T}$ is large and dynamic oracle $\sum_{t=1}^{T}\max_{a\in\mathcal{A}}u^{t}(a)$ is small.
Fundamental limit
Finally, we also complement our upper bounds by matching lower bounds showing that $\mathrm{SAOL}^{\mathcal{H}}$ and behavior cloning are already the near-optimal algorithms in terms of rates when competing with dynamic oracles and zero payoff respectively. The techniques are based on adapting existing hard instances for a more general setup to the symmetric zero-sum game setting. Please see more discussion in Appendix D.
@@
There exists some absolute constant $C>0$ such that for any $n\geq 3$, $A\geq 2$, and $V_{T}\in[1,T]$, it holds that $\inf_{{\text{Alg}}}\sup_{\mathcal{G}(n,A,V_{T})}\mathbb{E}_{{\text{Alg}}}[{\text{D-Reg}(T)}]\geq CV^{1/3}_{T}T^{2/3},$ where the expectation $\mathbb{E}_{{\text{Alg}}}[\cdot]$ is taken over the noisy actions and the intrinsic randomness in the algorithm.
@@
There exists some absolute constant $C>0$ such that for any $n\geq 3$, $A\geq 2$, and $V_{T}\in[1,T]$, it holds that $\sup_{{\text{Alg}}}\inf_{\mathcal{G}(n,A,V_{T})}\mathbb{E}_{{\text{Alg}}}\left[\sum^{T}_{t=1}u^{t}(x^{t})\right]\leq-CV_{T},$ where $\{x^{t}\}^{T}_{t=1}$ is the policy sequence implemented by the algorithm, and the expectation $\mathbb{E}_{{\text{Alg}}}[\cdot]$ is taken over the noisy actions and the intrinsic randomness in the algorithm.

### v2

@@ 3 Insufficiency of Equilibria and Self-play for Equal Share
Every player chooses either 00 or $1$. If all players take the same action, then they receive a payoff of 0. Otherwise, being in the majority yields a positive payoff of $1/2$, while being in the minority results in a negative payoff of $-1$.
Insufficiency of equilibria
@@
In this setup, both pure strategies $(0,0,0)$ and $(1,1,1)$ constitute NE. However, the existence of multiple NEs creates a predicament for the learning agent. It must choose which equilibrium to follow, yet there is always the risk that the two opponents are both playing the other NE, leading to a negative payoff for the learner. In other words, adhering to a single NE does not reliably ensure an equal share when multiple equilibria exist. Since ${\rm NE}\subset{\rm CE}\subset{\rm CCE}$, we know the same limitation also holds for CE and CCE.
Insufficiency of self-play from scratch
Self-play is a training method in which the learning agent improves its performance by repeatedly playing against copies of itself without human supervision. See pseudo-code in Algorithm 1. The learner maintains its own strategy $\{x_{t}\}_{t=1}^{T}$. At the $t^{\text{th}}$ iteration, the learner first pretends that all opponents are employing its current strategy $x_{t}$, and samples actions from them. Then the learner updates its own strategy to $x_{t+1}$ using the gradient information from the gameplay. The updates can be made using any optimizer such as gradient descent or Hedge.
Algorithm 1 Self-play meta-algorithm
1:  Initialize learner’s mixed strategy $x_{1}$.
2:  for $t=1,\ldots,T$ do
3:     Sample action $a_{i}^{t}\sim x_{t}$ for all player $i\in[n]$.
4:     Update strategy $x_{t+1}$ using the gradient information $U_{1}(\cdot,a_{-1}^{t})$.
5:  end for
Here, we argue that self-play from scratch (the algorithm adopted in Brown & Sandholm (2019)) again fails to secure an equal share in the same three-player majority vote game: Consider two opponents play the same fixed strategy that is one of the NEs. In this case, the learner has no choice but to play the exact same NE as the opponents to secure an equal share. That is, if the learner’s algorithm is agnostic to the strategies of the opponents, it is doomed to fail. We note that while recent systems (Li et al., 2020; Jacob et al., 2022) combine self-play with behavior cloning which is no longer agnostic to opponents’ strategies, our experiment shows that their meta-algorithms remain insufficient to secure an equal share in the worst-case scenario (See Section 6).
4 Sufficient Conditions for Securing Equal Share
In this section, we identify the structural conditions of the games where equal share is achievable. We will show that the following two conditions are needed to achieve equal share:
  1. Condition 1.
    All opponents need to deploy the same strategy, i.e., $x_{2}=\ldots=x_{n}$;
  2. Condition 2.
    All opponents must have limited adaptivity and the player has to model the opponents, i.e., $\{x_{j}\}_{j=2}^{n}$ can not adversarially change across different rounds of the game.
We justify these two conditions by proving that without either condition, an equal share can not be attained in multiplayer symmetric games in the worst case. We remark that both conditions restrict the strategies of opponents rather than the type of games.
Non-adaptive opponents that deploy different strategies
We start by considering the case where Condition 2 holds but Condition 1 does not hold.
@@ 4 Sufficient Conditions for Securing Equal Share
There exist symmetric zero-sum games with opponents using fixed but differing strategies, such that no learner’s strategy secures an equal share. In math,
$\max_{x_{1}}\min_{x_{2},\cdots,x_{n}}U_{1}(x_{1},\cdots,x_{n})\leq\min_{x_{2},\cdots,x_{n}}\max_{x_{1}}U_{1}(x_{1},\cdots,x_{n})\leq 0$ (1)
where both inequalities can be made strict in certain games.
Here, we can further strengthen the proposition to require opponents to deploy strategies without “collusion”. That is, the hard instance holds even when the strategies employed by opponents are statistically independent without any shared randomness. Proposition 4.1 highlights the challenge when opponents are free to adopt different strategies.
Adaptive opponents that deploy identical strategy
We next examine the case where Condition 1 holds but Condition 2 does not hold.
@@
There exist symmetric zero-sum games such that no learner’s strategy secures an equal share against adversarial opponents, even under the constraint that they adhere to identical strategy at each round. In math,
$\max_{x_{1}}\min_{x}U_{1}(x_{1},x^{\otimes n-1})\leq\min_{x}\max_{x_{1}}U_{1}(x_{1},x^{\otimes n-1})=0,$ (2)
where the inequality can be made strict in certain games.
Proposition 4.2 implies a property that makes multiplayer games significantly different from two-player zero-sum games: even under the favorable scenario of all opponents employing identical strategies, one can no longer find a fixed “non-exploitable” strategy agnostic to the strategies of opponents. All strategies are exploitable. Opponent modeling is necessary to secure an equal share.
Solution concepts beyond equilibrium
Combining Proposition 4.1 and Proposition 4.2, we observe that both conditions mentioned at the beginning of Section 4 are needed to make equal share achievable. We conclude from (1) and (2) that, out of the four related minimax concepts, only $\min_{x}\max_{x_{1}}U_{1}(x_{1},x^{\otimes n-1})=0$ across all multiplayer symmetric zero-sum games, which guarantees an equal share. This remaining minimax concept precisely corresponds to the two identified conditions where opponents employ identical strategies, and the learner must be adaptive to the opponents. Therefore, we will use $\min_{x}\max_{x_{1}}U_{1}(x_{1},x^{\otimes n-1})$ as our target solution concept for this paper to achieve equal share. We remark that this solution concept does not necessarily correspond to any equilibrium in most multiplayer games. We conclude with this solution concept from a principled manner with equal share as our primary objective. For conciseness, from now on, we will also refer to the common strategy employed by all opponents as the meta-strategy.
4.1 Connections between identified conditions and practice
While the two identified conditions may seem restrictive, here we argue that they in fact apply to practical multiplayer gaming platforms with a large player base. Condition 1 is further implicitly adopted by most prior state-of-the-art AI agents for multiplayer games.
Connection to multiplayer games with a large player base
We argue that both identified conditions are well-justified in modern multiplayer gaming platforms with a large player base. Imagine a casino hosting $N$ players who randomly join poker tables or an online Mahjong match-making platform with $N$ users. Let $\{x_{i}\}^{N}_{i=1}$ be the strategy set for these $N$ players. We can then define the population meta-strategy as $\bar{x}=(1/N)\sum_{i=1}^{N}x_{i}$. The following proposition claims that, for $n$-player symmetric zero-sum games, in terms of the expected payoff, playing against $n-1$ random players is almost the same as playing against $n-1$ players who all adopt the same population meta-strategy $\bar{x}$, as long as $N\gg(n-2)^{2}$.
Proposition 4.3.
Let $\mathbb{E}_{x_{-1}}$ be the expectation over the randomness on sampling $n-1$ strategies uniformly from the set $\{x_{i}\}^{N}_{i=1}$ without replacement. Then for any strategy $z\in\Delta(\mathcal{A})$, we have $|\mathbb{E}_{x_{-1}}[U_{1}(z,x_{-1})]-U_{1}(z,\bar{x}^{\otimes n-1})|\leq 2(n-2)^{2}/N$.
Furthermore, it is often safe to assume that the population meta-strategy $\bar{x}$ — the average strategy of all players within the player pool — will not quickly adapt to one particular player’s strategy.
Connection to practical AI systems
We remark that a majority of practical AI systems for multiplayer games (Li et al., 2020; Brown & Sandholm, 2019; Bakhtin et al., 2022; † et al.(2022)(FAIR)†, Bakhtin, Brown, Dinan, Farina, Flaherty, Fried, Goff, Gray, Hu, et al., FAIR) leverage self-play meta-algorithms (Algorithm 1), which equate the strategies of all opponents with those of the learner. This implicitly assumes all opponents employ an identical strategy at every round.
5 Provably Efficient Algorithms
In this section, we explore efficient algorithms that provably secure an equal share under the two conditions identified in Section 4. Particularly, we consider several opponent settings with various adaptivity: fixed, slowly adapting, and opponents that adapt at intermediate rates.
We use the following notations throughout this section: at round $t$, let $x^{t}$ denote the learner’s strategy, and $y^{t}$ the meta-strategy employed by all opponents. We denote $u^{t}(\cdot)$ as the expected payoff function of the learner at round $t$, where $u^{t}(\cdot):=U_{1}(\cdot,(y^{t})^{\otimes n-1})$. Then the average payoff of the learner is:
${u_{\text{avg}}(T)}:=\textstyle(1/T)\sum^{T}_{t=1}u^{t}(x^{t}).$
@@ 5.1 Fixed opponents
We begin by exploring the simple stationary scenario, where the meta-strategy used by the opponents remains constant over time, denoted as $y^{t}=y$ for all $t\in[T]$.
Notably, in this particular scenario, the payoff function $u^{t}(\cdot)$ remains constant over time. Additionally, by symmetry, it is not hard to observe that at least one action will consistently yield an expected payoff of 0 in all rounds. This implies $\max_{a\in\mathcal{A}}\sum_{t=1}^{T}u^{t}(a)\geq 0$, which makes the no-regret learning tool well-poised to achieve equal share in this setting. Standard (static) regret, defined as follows, compares the learner’s total payoff to the total payoff achieved by the best action in hindsight:
${\text{Reg}(T)}:=\max_{a\in\mathcal{A}}\sum_{t=1}^{T}u^{t}(a)-\sum^{T}_{t=1}u^{t}(x^{t}),$
An algorithm has no-regret if ${\text{Reg}(T)}\leq o(T)$ for all large $T$. We deploy a standard no-regret learning algorithm — Hedge (Freund & Schapire, 1997), which provides the following guarantees:
@@
${u_{\text{avg}}(T)}\geq u^{\star}-C\sqrt{\log(A/\delta)/T},$
for some absolute constant $C$, where $u^{\star}:=\max_{a\in\mathcal{A}}U_{1}(a,y^{\otimes n-1})\geq 0$.
The probability is taken over the random actions by opponents sampled from the meta-strategy [^5] Recall that our learning protocol in Section 2 assumes the learner only observes the noisy actions of the other players at each round.. Theorem 5.1 claims that with stationary opponents, the Hedge algorithm approximately achieves an equal share up to a $\tilde{O}(1/\sqrt{T})$ error, which demonstrates its effectiveness in the long run.
@@ 5.2 Adaptive opponents
In practical scenarios, encountering a fixed opponent strategy is relatively uncommon. More often, opponents adapt and modify their strategies over time, responding to the game’s dynamics and the actions of other players. Thus, in this section, we shift our focus to the non-stationary scenario, where the meta-strategy $y^{t}$ adopted by the opponents varies over time.
According to Proposition 4.2, it is clear that attaining an equal share is impossible if opponents can change their meta-strategy $y^{t}$ arbitrarily fast across different rounds. Thus we introduce a constraint on the adaptive power of the opponents by positing a variation budget $V_{T}$, which bounds the total variation of the payoff function across the time horizon. Specifically, we assume the payoff function belongs to $\mathcal{U}$, which is defined as
$\mathcal{U}:=\left\{\{u^{t}\}^{T}_{t=1}\,\left|\,\sum_{t=1}^{T-1}\left\|u^{t+1}-u^{t}\right\|_{\infty}\leq V_{T}\right.\right\}.$ (3)
Furthermore, we denote $\mathcal{G}(n,A,V_{T})$ as the set of tuples, which consists of a $n$-player symmetric zero-sum game with $A$ actions and a corresponding meta-strategy sequence $\{y^{t}\}_{t=1}^{T}$, such that the payoff function $\{u^{t}\}^{T}_{t=1}\in\mathcal{U}$. This constraint effectively moderates the adaptivity of the opponents compared to a fully adversarial setup.
@@
In non-stationary environments, the total payoff achieved by the best action in hindsight $\max_{a\in\mathcal{A}}\sum_{t=1}^{T}u^{t}(a)$ is no longer non-negative. Therefore, minimizing standard regret in this setting is no longer effective in securing an equal share. This motivates us to turn our attention to a stronger notion of regret — dynamic regret, defined as:
@@
In the setting of symmetric games, the dynamic oracle is always assured to secure an equal share, i.e., $\sum_{t=1}^{T}\max_{a\in\mathcal{A}}u^{t}(a)\geq 0$. Thus, any algorithm achieving no-dynamic-regret is guaranteed to achieve an equal share up to a small error. To this ends, we adapt a no-dynamic-regret algorithm—Strongly Adaptive Online Learner with Hedge $\mathcal{H}$ as a black box algorithm ($\mathrm{SAOL}^{\mathcal{H}}$) (See Appendix C.1.2), as proposed by Daniely et al. (2015) to our setting and achieve following guarantees:
@@
Suppose that $n\geq 3$, $A\geq 2$, and $V_{T}\in[1,T]$, then for any game and any meta-strategy sequence in $\mathcal{G}(n,A,V_{T})$, with probability at least $1-\delta$, $\mathrm{SAOL}^{\mathcal{H}}$ satisfies
${u_{\text{avg}}(T)}\geq u^{\dagger}-CV^{1/3}_{T}T^{-1/3}\left(\sqrt{\log(A/\delta)}+\log T\right)$
for some absolute constant $C$, where $u^{\dagger}:=(1/T)\sum_{t=1}^{T}\max_{a\in\mathcal{A}}u^{t}(a)\geq 0$.
Theorem 5.2 implies that $\mathrm{SAOL}^{\mathcal{H}}$ achieves a non-negative average payoff, up to an error term that scales with $\tilde{O}(V_{T}^{1/3}T^{-1/3})$. Therefore, if $V_{T}$ is sublinear in $T$, $\mathrm{SAOL}^{\mathcal{H}}$ is capable of approximately achieving equal share over an extended duration.
Opponents that adapt at intermediate rates
Interestingly, there is an intermediate regime where opponents’ strategies $\{y^{t}\}_{t=1}^{T}$ are changing neither too fast nor too slow where the favorable algorithm for the learner might be simply behavior cloning—simply mimic opponents’ strategies.
Formally, we define the behavior cloning algorithm for the learner by making her action in $t$-th round the same as the action taken by the 2nd player in $(t-1)$-th round (See Algorithm 2). Behavior cloning achieves the following:
@@
Suppose that $n\geq 3$, $A\geq 2$, and $V_{T}\in[1,T]$, for any game and any meta-strategy sequence in $\mathcal{G}(n,A,V_{T})$, behavior cloning guarantees that
$\mathbb{E}[{u_{\text{avg}}(T)}]\geq-(V_{T}+1)/T.$
We remark that while the error term $O(V_{T}/T)$ in Theorem 5.3 is always smaller than the error term $\tilde{O}((V_{T}/T)^{1/3})$ in Theorem 5.2, the latter is comparing to the baseline of dynamic oracle $u^{\dagger}$, which has a greater value than the baseline in behavior cloning — 00. It is not hard to see that in the intermediate regime $\tilde{\Theta}(u^{\dagger})\leq V_{T}/T\leq\Theta(1)$, the meta-strategy is changing too fast so that it is better for the learner to simply copy the meta-strategy instead of running sophisticated no-dynamic-regret algorithm to learn the game and to counter the meta-strategy by herself.
Matching lower bounds
Finally, we also complement our upper bounds by matching lower bounds showing that $\mathrm{SAOL}^{\mathcal{H}}$ and behavior cloning are already the near-optimal algorithms in terms of error rates when compared with the corresponding baselines — the dynamic oracle and zero payoff respectively. The techniques used here are based on adapting existing hard instances for a more general setup to the symmetric zero-sum game setting. Please see more discussion in Appendix C.2.
@@
There exists some absolute constant $C>0$ such that for any $n\geq 3$, $A\geq 2$, and $V_{T}\in[1,T]$, and any learning algorithm, there exists a game and a meta-strategy sequence in $\mathcal{G}(n,A,V_{T})$, such that $\mathbb{E}[{u_{\text{avg}}(T)}]\leq u^{\dagger}-CV^{1/3}_{T}T^{-1/3}$.
@@
There exists some absolute constant $C>0$ such that for any $n\geq 3$, $A\geq 2$, and $V_{T}\in[1,T]$, and any learning algorithm, there exists a game and a meta-strategy sequence in $\mathcal{G}(n,A,V_{T})$ such that $\mathbb{E}\left[{u_{\text{avg}}(T)}\right]\leq-CV_{T}/T$.
The expectation in both theorems are taking over the random actions by the opponents as well as the possible intrinsic randomness in a stochastic algorithm. Theorem 5.4 and 5.5 match Theorem 5.2 and 5.3 respectively, up to additional logarithmic factors.